### v1

# Evaluating GPT-V4 (GPT-4 with Vision) on Detection of Radiological Findings on Chest Radiographs

Abstract
Background
Generating radiological findings from chest radiographs is pivotal in medical image analysis. The emergence of GPT-4 with vision (GPT-4V) has opened new perspectives on the potential for automated image-text pair generation. However, the application of GPT-4V to real-world chest radiography is yet to be thoroughly examined.
Purpose
To investigate GPT-4V’s capability to generate radiological findings from real-world chest radiographs.
Materials and Methods
In this retrospective study, 100 chest radiographs with free-text radiology reports were annotated by a cohort of two radiology-attending physicians and three residents to establish a reference standard. Out of 100 chest radiographs, 50 were randomly picked from the National Institutes of Health (NIH) chest radiography data set, and 50 were randomly selected from the Medical Imaging and Data Resource Center (MIDRC). The performance of GPT-4V at detecting imaging findings from each chest radiograph was assessed in the zero-shot settings (where it operates without prior examples) and few-shot settings (where it operates with two examples). Its outcomes were compared with the reference standard about radiological conditions, their corresponding codes in the International Classification of Diseases, Tenth Revision (ICD-10 codes), and laterality.
Results
In the zero-shot setting, in the task of detecting ICD-10 codes alone, GPT-4V attained a G&R+/R+ of 12.3%, a G&R+/G+ of 5.8%, and an F1 score of 7.3% on the NIH dataset, and a G&R+/R+ of 25.0%, a G&R+/G+ of 16.8%, and an F1 score of 18.2% on the MIDRC dataset. When both the ICD-10 codes and their corresponding laterality were considered, GPT-4V produced a G&R+/R+ of 7.8%, a G&R+/G+ of 3.5%, and an F1 score of 4.5% on the NIH dataset, and a G&R+/R+ of 3.6%, a G&R+/G+ of 4.9%, and an F1 of 6.4% on the MIDRC dataset. In few-shot learning, GPT-4V showed improved performance on both datasets. When contrasting zero-shot and few-shot learning, there were improved G&R+/G+ and F1 scores in few-shot learning. Nonetheless, there was not a substantial increase in the G&R+/R+.
Conclusion
Although GPT-4V showed promise in understanding real-world images but had limited effectiveness in interpreting real-world chest radiographs.
1 Introduction
Generating radiological findings from chest radiographs is pivotal in medical image analysis[1]. Recent advancements in fine-tuned pretrained models have showcased their capability to translate image content into text[2]. However, these models are often trained on extensive, non-specific datasets and may need more domain-specific tuning for chest radiographs. The emergence of OpenAI’s GPT-4 with vision (GPT-4V), a multimodal large language model (LLM) with visual recognition, has opened new perspectives on the potential for automated image-text pair generation in the medical care domain. Advanced multimodal LLMs, such as GPT-4V, can understand both text and images. While several studies have investigated the performance of GPT-4 in generating radiological impressions[3] and summarizing clinical trials[4], the practical application of multimodal LLMs to real-world chest radiographs is yet to be thoroughly examined. Motivated by this gap, this study aimed to investigate GPT-4V’s capability to generate radiological findings from real-world chest radiographs.
2 Materials & Methods
Because of the publicly available nature of the data set used in this study, the requirement to obtain written informed consent from all subjects was waived by the institutional review board.
2.1 Study Design and Data Collection
In this retrospective study, a total of 100 chest radiographs and reports were independently annotated by a cohort of two radiology-attending physicians and three residents to establish a reference standard (Figure 1). Out of 100 chest radiographs, 50 chest radiographs were randomly picked from the National Institutes of Health (NIH) chest radiography data set, and their corresponding reports were dictated by one radiology-attending physician and three radiology residents[3]. These 50 patients have been previously reported[4]. The prior article dealt with the generation of the impression section using the finding section in the reports, whereas this article reports on the generation of a radiological finding table from the images.
The remaining 50 chest radiographs and de-identified free-text radiology reports were randomly selected from the Medical Imaging and Data Resource Center (MIDRC)[4]. Each report included a Findings section and an Impressions section.
Out of 100 chest radiographs and reports, 10 cases were randomly selected (5 from the NIH dataset and 5 from the MIDRC dataset), and two were randomly selected to serve as few-shot examples for the GPT-4V model. The remaining 90 were used to evaluate GPT-4V’s performance in zero-shot learning scenarios, where it operates without prior examples, and in few-shot learning scenarios, where it operates with two examples. Its outcomes were then compared with the reference standard annotated by radiologists about clinical conditions, their corresponding codes in the International Classification of Diseases, Tenth Revision (ICD-10), and laterality.
Figure 1: The workflow of data construction and model application. GPT-4V=GPT-4 with vision, MIDRC= Medical Imaging and Data Resource Center, NIH= National Institutes of Health
2.2 GPT-4 with Vision (GPT-4V)
GPT-4V (Oct.13th, 2023 version; OpenAI) was used. This version of GPT enables users to instruct GPT-4 to analyze image inputs.
2.3 Experimental setup
To obtain the reference standard tables, GPT-4 was used to convert each free-text radiology report (Figure 2a) into a table of radiological findings with the prompt shown in Figure 2b. This table included the radiological findings, the corresponding ICD-10 diagnostic codes and their laterality, and detailed descriptions of the ICD codes (Figure 2c). Subsequently, each report was independently evaluated by three readers from a cohort of five board-certified radiologists and residents (H.O., P.K., C.W., J.K., G.S). Two of the readers are third-year radiology residents, and the remaining three are radiology attendings, each with over 15 years of experience. Their radiology subspecialties cover chest, ED/bone, neurology, and body imaging. All readers had access to the image views and reports but not to additional clinical or patient data. Both datasets, the National Institutes of Health (NIH) and the Medical Imaging and Data Resource Center (MIDRC), were comprehensively reviewed, with each dataset’s 50 reports being examined by three reviewers to maintain consistency and objectivity in the evaluation process. Findings were only included in the final tables if they were observed by at least two of the three reviewers; conversely, findings were excluded if two or more reviewers did not identify them. The majority vote principle was employed to provide a clear consensus for the presence or absence of radiological findings and lead to the final reference-standard table.
Figure 2: An example. a, The radiology report. b, The prompt used to convert the free-text report into a table of radiological findings. c, The table of radiological findings that has been confirmed by radiologists using the report. d, The chest radiography input of GPT-V. e, The prompt utilized by GPT-4V to create a table of radiological findings derived from the chest radiography. f, The result produced by GPT-4V.
2.3.1 Evaluating Performance in the zero-shot setting
To evaluate GPT-4V’s performance in the zero-shot setting, the chest radiographs (Figure 2d) were inputted into the GPT-4V model (version from Oct.13th)[2] along with the prompt (Figure 2e). This step aims to generate a radiological findings table (Figure 2f) comparable to the reference standard. The analysis concentrated on aligning positive radiological finding identification and laterality between GPT-generated and consensus tables. G&R+/R+ (positive predictive value), G&R+/G+ (true positive rate), and F1 scores were calculated at the report level (detailed in the section of “Evaluation metrics”). These metrics were used to assess the accuracy of GPT-4V in detecting the ICD-10 codes and their respective laterality.
2.3.2 Evaluating Performance in the few-shot setting
To evaluate GPT-4V’s performance in the few-shot setting, the input was extended to include two examples of chest radiographs with their corresponding radiological findings tables before the prompt. From the pool of 10 chest radiographs, two were randomly selected to serve as few-shot examples for the GPT-4V model. Supplying the model with these examples helped to provide context, boosting the model’s capacity to generate an accurate radiological findings table. The same performance metrics— G&R+/R+, G&R+/G+, and F1 score— were used at the report level to assess the effectiveness of few-shot learning.
2.4 Statistical Analysis
When obtaining the final reference standard tables, the interrater agreement was reported using Cohen’s Kappa score[5].
First, the GPT-4V was employed to detect radiological findings from each chest radiography and compared the predicted findings with those obtained from the radiologists. To convey the performance evaluation, G&R+ (true positive rate) was used to represent the number of ICD-10 Codes correctly predicted by GPT-4V, R+ for the total number of ICD-10 Codes identified by GPT-4V, and G+ for the count of ICD-10 Codes existing in the reference standard.
Evaluation metrics of G&R+/R+ (positive predictive value, PPV), G&R+/G+ (true positive rate, TPR), and F1 score were used to assess GPT-4V’s performance in detecting imaging findings from each chest radiograph [6]. These predicted findings were then compared with those obtained from the radiologists.
$F1=2*\frac{PPV*TPR}{PPV+TPR}$
In the evaluation of detecting ICD-10 codes, a radiological finding was considered a true positive if its ICD-10 code aligned with that in the reference standard table. In evaluating both the radiological findings in ICD codes and their corresponding lateralities, a radiological finding was considered a true positive if both its ICD-10 code and laterality matched those in the reference standard table.
For example, in Figure 2, while evaluating ICD-10 codes alone, there were 2 G&R+ (J90 and J18.9), five findings predicted by GPT-4V, and four in the reference standard table. Therefore, G&R+/R+ was 0.4 (=2/(2+3)) , G&R+/G+ was 0.5 (=2/(2+2)), and F1 score was 0.44.
After obtaining the G&R+/R+, G&R+/G+, and F1 score for each chest radiography, the macro averages were calculated for G&R+/R+, G&R+/G+, and F1 score across all 90 chest radiographs.
In this study, the statistical significance was set at 0.05, and two-tailed t-tests were used for calculating the P values for the performance metrics of GPT-4V in detecting radiological findings from chest radiography, specifically assessing the detection of findings represented by ICD codes and their associated lateralities in both zero-shot and few-shot settings.
3 Results
3.1 Performance in the zero-shot setting
The performance of the GPT-4V model in the zero-shot setting varied across the NIH and MIDRC datasets and under different circumstances (Figure 3). In the task of detecting ICD-10 codes alone, the model attained a G&R+/R+ of 5.53/45 (12.3%), with a standard deviation (SD) of 0.25, a standard error (SE) of 0.04, and an interquartile range (IQR) of 0.20; a G&R+/G+ of 2.60/45 (5.8%), with a SD of 0.10, a SE of 0.02, and an IQR of 0.10; and an F1 score of 3.30/45 (7.3%), with a SD of 0.13, a SE of 0.02, and an IQR of 0.14, on the NIH dataset. Conversely, on the MIDRC dataset, the model managed a G&R+/R+ of 11.25/45 (25.0%), with a SD of 0.24, a SE of 0.04, and an IQR of 0.33; a G&R+/G+ of 7.56/45 (16.8%), with a SD of 0.20, a SE of 0.03, and an IQR of 0.25; and an F1 score of 8.20/45 (18.2%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.29. The notable differences in measurements between the two datasets were primarily due to the MIDRC dataset having fewer missing GPT-4V-generated ICD-10 codes than the NIH dataset. On the MIDRC dataset, GPT-4V generated 144 radiological findings, while the reference standard comprised 261 findings. However, on the NIH dataset, GPT-4V produced 102 radiological findings, while the reference standard comprised 220 findings. Nevertheless, when both the ICD-10 codes and their corresponding laterality were taken into account, the GPT-4V model in the zero-shot setting produced a G&R+/R+ of 3.5/45 (7.8%), with a SD of 0.20, a SE of 0.03, and an IQR of 0.0; a G&R+/G+ of 1.56/45 (3.5%), with a SD of 0.01, a SE of 0.01, and an IQR of 0.0; and an F1 score of 2.05/45 (4.5%) with a SD of 0.10, a SE of 0.02, and an IQR of 0.0 on the NIH dataset; and a G&R+/R+ of 4.92/45 (3.6%), with a SD of 0.21, a SE of 0.03, and an IQR of 0.25; a G&R+/G+ of 2.19/45 (4.9%), with a SD of 0.09, a SE of 0.01, and an IQR of 0.10; and an F1 of 2.90/45 (6.4%), with a SD of 0.11, a SE of 0.02, and an IQR of 0.15, on the MIDRC dataset.
Figure 3: Performance of GPT-4 with Vision (GPT-4V) in the detection of radiological findings from chest radiography in the zero-shot setting, with statistical significance assessed using a two-tailed t-test. a, Detecting the radiological findings in ICD codes. b, Detecting both the radiological findings in ICD codes and their corresponding lateralities. * - p < 0.05; ** - p < 0.01; *** - p < 0.001; ns – Not significant. ICD= International Classification of Diseases, Tenth Revision.
3.2 Performance in the few-shot setting
In few-shot learning, GPT-4V showed improved performance on both NIH and MIDRC datasets (Figure 4). When the model was provided with two illustrative chest radiographs and reports, there was a marked improvement in the G&R+/R+ on the NIH dataset, with rates rising to 5.72/45 (12.7%) with a SD of 0.19, a SE of 0.03, and an IQR of 0.25. G&R+/G+ also enhanced to 4.69/45 (10.4%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.22; and the F1 score reached 5.03/45 (11.1%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.22. On the MIDRC dataset, the G&R+/R+ was improved to 16.15/45 (35.9%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50; the G&R+/G+ improved to 16.68/45 (37.1%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50; and the F1 score improved to 15.47/45 (34.3%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50. When tasked with detecting both ICD-10 codes and their corresponding lateralities, the GPT-4V model demonstrated improved efficacy. On the NIH dataset, it displayed a G&R+/R+ of 1.62/45(3.5%), with a SD of 0.10, a SE of 0.01, and an IQR of 0.0; a G&R+/G+ of 1.14/45 (2.5%), with a SD of 0.07, a SE of 0.01, and an IQR of 0.0; and an F1 score of 1.30/45(2.8%), with a SD of 0.08, a SE of 0.01, and an IQR of 0.0. On the MIDRC dataset, it achieved a G&R+/R+ of 8.96/45 (19.9%), with a SD of 0.22, a SE of 0.03, and an IQR of 0.33; a G&R+/G+ of 9.14/45 (20.3%), with a SD of 0.25, a SE of 0.04, and an IQR of 0.33; and an F1 score of 8.53/45 (19.0%), with a SD of 0.21, a SE of 0.03, and an IQR of 0.31.
Figure 4: Performance of GPT-4 with Vision (GPT-4V) in the detection of radiological findings from chest radiography in the few-shot setting, with statistical significance assessed using two-tailed t-test. a, Detecting the radiological findings in ICD codes. b, Detecting both the radiological findings in ICD codes and their corresponding lateralities. * - p < 0.05; ** - p < 0.01; *** - p < 0.001; **** - p < 0.0001; ns – Not significant. ICD= International Classification of Diseases, Tenth Revision.
When contrasting zero-shot and few-shot learning approaches on both datasets (Figure 5), there were improved G&R+/G+ and F1 scores in few-shot learning in both scenarios (ICD-10 codes only and ICD-10 codes with laterality). Nonetheless, there was not a substantial increase in the G&R+/R+ suggesting that while few-shot learning may enhance the model’s capacity to detect findings, it does not noticeably enhance the precision of the predictions.
Figure 5: Difference of performance of GPT-4 with Vision (GPT-V4) in the detection of radiological findings from chest radiography between zero-shot and few-shot settings, with statistical significance assessed using two-tailed t-test. a, Detecting the radiological findings in ICD codes. b, Detecting both the radiological findings in ICD codes and their corresponding lateralities.
3.3 Inter-rater Agreement
The study compared each reader’s outcomes to the reference standard to evaluate their interpretations (Table 1). On the NIH dataset, the IRA for Reviewer A was 0.78, for Reviewer C was -0.06, and for Reviewer D was 0.96. On the MIDRC dataset, the IRA for Reviewer B was 0.96, for Reviewer C was 0.82, and for Reviewer D was 0.99.
Table 1: Reviewer Agreement (Cohen’s Kappa) for NIH and MIDRC. The Inter-rater Agreement (IRA) between each reviewer and the reference standard table for both NIH and MIDRC datasets. $k=1$ indicates perfect agreement among reviewers and the reference standard; $k=0$ indicates that agreement is no better than chance; $k=\ -1$ indicates perfect disagreement.
NIH Results MIDRC Results
Reviewer A 0.78 NA
Reviewer B NA 0.96
Reviewer C -0.06 0.82
Reviewer D 0.96 0.99
4 Discussion
The emergence of multimodal large language models (LLMs) that can understand both text and images, such as GPT-4 with Vision (GPT-4V), shows potential for automated image-text pair generation. However, applying these models to real-world data is yet to be thoroughly examined. This study assessed the feasibility of using GPT-4V to detect radiological findings from chest radiographs in both zero-shot and few-shot learning contexts. The results (G&R+/R+ of 5.53/45 (12.3%) and G&R+/G+ of 2.60/45 (5.8%)) demonstrated that radiological findings tables generated by the GPT-4V model still need to be prepared for use in clinical practice. We acknowledged a limitation in employing GPT-4 for converting radiological reports into a structured table, where inaccuracies in ICD-10 code assignments and distinguishing between radiological findings and conclusions may impact the data’s reliability and interpretability. A notable limitation of the GPT-4V output was its failure to detect several radiological conditions based on corresponding codes in the International Classification of Diseases, Tenth Revision (ICD-10). Overall, the top three findings that GPT-4V could not detect were “endotracheal tube”, “central venous catheter”, and “degenerative changes of osseous structures”. Conversely, GPT-4V most accurately detected findings such as “chest drain”, “air-space disease”, and “lung opacity”.
Although GPT-4V showed promise in understanding real-world images[7], its effectiveness in interpreting real-world chest radiographs was limited.
Our study had limitations. First, the few-shot training is that it may be more prone to generating ICD-10 codes already in the provided examples, potentially reducing the diversity of ICD-10 codes generated. Second, we were unable to access other multimodal LLMs that support image inputs. Consequently, we lacked comparative data with respect to the results of GPT-4v. Finally, due to the limited size of the dataset and relatively low level of inter-rater agreement among the radiologists, analysis by GPT-4V may have been challenging.
In conclusion, GPT-4 with Vision (GPT-4V) showed promise in understanding natural images but had limited effectiveness in interpreting real-world chest radiographs. Our results highlight the need for additional comprehensive development and assessments prior to incorporating the GPT-4V model into clinical practice routines. Task-specific, fine-tuned multimodal LLMs or foundation models are urgently needed for this purpose, although it is not necessarily the best solution. To yield robust and generalizable results, we plan to explore larger and more diverse datasets using real-world data in future studies. This will involve including multiple modalities, such as brain CT scans and MRIs, to conduct a more thorough evaluation of GPT-4V’s performance.

### v2

@@
 Results
-In the zero-shot setting, in the task of detecting ICD-10 codes alone, GPT-4V attained a G&R+/R+ of 12.3%, a G&R+/G+ of 5.8%, and an F1 score of 7.3% on the NIH dataset, and a G&R+/R+ of 25.0%, a G&R+/G+ of 16.8%, and an F1 score of 18.2% on the MIDRC dataset.
-When both the ICD-10 codes and their corresponding laterality were considered, GPT-4V produced a G&R+/R+ of 7.8%, a G&R+/G+ of 3.5%, and an F1 score of 4.5% on the NIH dataset, and a G&R+/R+ of 3.6%, a G&R+/G+ of 4.9%, and an F1 of 6.4% on the MIDRC dataset.
+In the zero-shot setting, in the task of detecting ICD-10 codes alone, GPT-4V attained an average PPV of 12.3%, an average TPR of 5.8%, and an average F1 score of 7.3% on the NIH dataset, and an average PPV of 25.0%, an average TPR of 16.8%, and an average F1 score of 18.2% on the MIDRC dataset.
+When both the ICD-10 codes and their corresponding laterality were considered, GPT-4V produced an average PPV of 7.8%, an average TPR of 3.5%, and an average F1 score of 4.5% on the NIH dataset, and an average PPV of 3.6%, an average TPR of 4.9%, and an average F1 of 6.4% on the MIDRC dataset.
 In few-shot learning, GPT-4V showed improved performance on both datasets.
-When contrasting zero-shot and few-shot learning, there were improved G&R+/G+ and F1 scores in few-shot learning.
-Nonetheless, there was not a substantial increase in the G&R+/R+.
+When contrasting zero-shot and few-shot learning, there were improved average TPR and F1 scores in few-shot learning.
+Nonetheless, there was not a substantial increase in the average PPV.
 Conclusion
@@
 Subsequently, each report was independently evaluated by three readers from a cohort of five board-certified radiologists and residents (H.O., P.K., C.W., J.K., G.S).
-Two of the readers are third-year radiology residents, and the remaining three are radiology attendings, each with over 15 years of experience.
+Two of the readers were third-year radiology residents, and the remaining three are radiology attendings, each with over 15 years of experience.
 Their radiology subspecialties cover chest, ED/bone, neurology, and body imaging.
@@
 2.3.1 Evaluating Performance in the zero-shot setting
-To evaluate GPT-4V’s performance in the zero-shot setting, the chest radiographs (Figure 2d) were inputted into the GPT-4V model (version from Oct.13th)[2] along with the prompt (Figure 2e).
+To evaluate GPT-4V’s performance in the zero-shot setting, the chest radiographs (Figure 2d) were input into the GPT-4V model (version from Oct.13th)[2] along with the prompt (Figure 2e).
 This step aims to generate a radiological findings table (Figure 2f) comparable to the reference standard.
 The analysis concentrated on aligning positive radiological finding identification and laterality between GPT-generated and consensus tables.
-G&R+/R+ (positive predictive value), G&R+/G+ (true positive rate), and F1 scores were calculated at the report level (detailed in the section of “Evaluation metrics”).
+Positive predictive value (PPV), true positive rate (TPR), and F1 scores were calculated at the report level (detailed in the section of “Evaluation metrics”).
+Notably, the application of the F1 scores represented an unconventional approach due to the potential involvement of multiple diagnoses.
 These metrics were used to assess the accuracy of GPT-4V in detecting the ICD-10 codes and their respective laterality.
@@
 Supplying the model with these examples helped to provide context, boosting the model’s capacity to generate an accurate radiological findings table.
-The same performance metrics— G&R+/R+, G&R+/G+, and F1 score— were used at the report level to assess the effectiveness of few-shot learning.
+The same performance metrics— PPV, TPR, and F1 score— were used at the report level to assess the effectiveness of few-shot learning.
 2.4 Statistical Analysis
@@
 First, the GPT-4V was employed to detect radiological findings from each chest radiography and compared the predicted findings with those obtained from the radiologists.
-To convey the performance evaluation, G&R+ (true positive rate) was used to represent the number of ICD-10 Codes correctly predicted by GPT-4V, R+ for the total number of ICD-10 Codes identified by GPT-4V, and G+ for the count of ICD-10 Codes existing in the reference standard.
-Evaluation metrics of G&R+/R+ (positive predictive value, PPV), G&R+/G+ (true positive rate, TPR), and F1 score were used to assess GPT-4V’s performance in detecting imaging findings from each chest radiograph [6].
+To convey the performance evaluation, positive predictive value (PPV) was utilized to denote the proportion of ICD-10 Codes correctly predicted by GPT-4V, while the true positive rate (TPR) represented the ration of true positive predictions to the total number of ICD-10 Codes identified by GPT-4V.
+Evaluation metrics of PPV, TPR, and F1 score were used to assess GPT-4V’s performance in detecting imaging findings from each chest radiograph [6].
 These predicted findings were then compared with those obtained from the radiologists.
@@
 In evaluating both the radiological findings in ICD codes and their corresponding lateralities, a radiological finding was considered a true positive if both its ICD-10 code and laterality matched those in the reference standard table.
-For example, in Figure 2, while evaluating ICD-10 codes alone, there were 2 G&R+ (J90 and J18.9), five findings predicted by GPT-4V, and four in the reference standard table.
-Therefore, G&R+/R+ was 0.4 (=2/(2+3)) , G&R+/G+ was 0.5 (=2/(2+2)), and F1 score was 0.44.
-After obtaining the G&R+/R+, G&R+/G+, and F1 score for each chest radiography, the macro averages were calculated for G&R+/R+, G&R+/G+, and F1 score across all 90 chest radiographs.
+For example, in Figure 2, while evaluating ICD-10 codes alone, there were 2 ICD-10 Codes correctly predicted by GPT-4V (J90 and J18.9), five findings predicted by GPT-4V, and four in the reference standard table.
+Therefore, PPV was 0.4 (=2/(2+3)) , TPR was 0.5 (=2/(2+2)), and F1 score was 0.44.
+After obtaining the PPV, TPR, and F1 score for each chest radiography, the macro averages were calculated for PPV, TPR, and F1 score across all 90 chest radiographs.
 In this study, the statistical significance was set at 0.05, and two-tailed t-tests were used for calculating the P values for the performance metrics of GPT-4V in detecting radiological findings from chest radiography, specifically assessing the detection of findings represented by ICD codes and their associated lateralities in both zero-shot and few-shot settings.
@@
 The performance of the GPT-4V model in the zero-shot setting varied across the NIH and MIDRC datasets and under different circumstances (Figure 3).
-In the task of detecting ICD-10 codes alone, the model attained a G&R+/R+ of 5.53/45 (12.3%), with a standard deviation (SD) of 0.25, a standard error (SE) of 0.04, and an interquartile range (IQR) of 0.20; a G&R+/G+ of 2.60/45 (5.8%), with a SD of 0.10, a SE of 0.02, and an IQR of 0.10; and an F1 score of 3.30/45 (7.3%), with a SD of 0.13, a SE of 0.02, and an IQR of 0.14, on the NIH dataset.
-Conversely, on the MIDRC dataset, the model managed a G&R+/R+ of 11.25/45 (25.0%), with a SD of 0.24, a SE of 0.04, and an IQR of 0.33; a G&R+/G+ of 7.56/45 (16.8%), with a SD of 0.20, a SE of 0.03, and an IQR of 0.25; and an F1 score of 8.20/45 (18.2%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.29.
+).
+In the task of detecting ICD-10 codes alone, the model attained an average PPV of 5.53/45 (12.3%), with a standard deviation (SD) of 0.25, a standard error (SE) of 0.04, and an interquartile range (IQR) of 0.20; an average TPR of 2.60/45 (5.8%), with a SD of 0.10, a SE of 0.02, and an IQR of 0.10; and an average F1 score of 3.30/45 (7.3%), with a SD of 0.13, a SE of 0.02, and an IQR of 0.14, on the NIH dataset.
+Conversely, on the MIDRC dataset, the model managed an average PPV of 11.25/45 (25.0%), with a SD of 0.24, a SE of 0.04, and an IQR of 0.33; an average TPR of 7.56/45 (16.8%), with a SD of 0.20, a SE of 0.03, and an IQR of 0.25; and an average F1 score of 8.20/45 (18.2%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.29.
 The notable differences in measurements between the two datasets were primarily due to the MIDRC dataset having fewer missing GPT-4V-generated ICD-10 codes than the NIH dataset.
@@
 However, on the NIH dataset, GPT-4V produced 102 radiological findings, while the reference standard comprised 220 findings.
-Nevertheless, when both the ICD-10 codes and their corresponding laterality were taken into account, the GPT-4V model in the zero-shot setting produced a G&R+/R+ of 3.5/45 (7.8%), with a SD of 0.20, a SE of 0.03, and an IQR of 0.0; a G&R+/G+ of 1.56/45 (3.5%), with a SD of 0.01, a SE of 0.01, and an IQR of 0.0; and an F1 score of 2.05/45 (4.5%) with a SD of 0.10, a SE of 0.02, and an IQR of 0.0 on the NIH dataset; and a G&R+/R+ of 4.92/45 (3.6%), with a SD of 0.21, a SE of 0.03, and an IQR of 0.25; a G&R+/G+ of 2.19/45 (4.9%), with a SD of 0.09, a SE of 0.01, and an IQR of 0.10; and an F1 of 2.90/45 (6.4%), with a SD of 0.11, a SE of 0.02, and an IQR of 0.15, on the MIDRC dataset.
+Nevertheless, when both the ICD-10 codes and their corresponding laterality were taken into account, the GPT-4V model in the zero-shot setting produced an average PPV of 3.5/45 (7.8%), with a SD of 0.20, a SE of 0.03, and an IQR of 0.0; an average TPR of 1.56/45 (3.5%), with a SD of 0.01, a SE of 0.01, and an IQR of 0.0; and an average F1 score of 2.05/45 (4.5%) with a SD of 0.10, a SE of 0.02, and an IQR of 0.0 on the NIH dataset; and an average PPV of 4.92/45 (3.6%), with a SD of 0.21, a SE of 0.03, and an IQR of 0.25; an average TPR of 2.19/45 (4.9%), with a SD of 0.09, a SE of 0.01, and an IQR of 0.10; and an average F1 of 2.90/45 (6.4%), with a SD of 0.11, a SE of 0.02, and an IQR of 0.15, on the MIDRC dataset.
 Figure 3: Performance of GPT-4 with Vision (GPT-4V) in the detection of radiological findings from chest radiography in the zero-shot setting, with statistical significance assessed using a two-tailed t-test.
@@
 b, Detecting both the radiological findings in ICD codes and their corresponding lateralities.
-* - p < 0.05; ** - p < 0.01; *** - p < 0.001; ns – Not significant.
 ICD= International Classification of Diseases, Tenth Revision.
@@
 In few-shot learning, GPT-4V showed improved performance on both NIH and MIDRC datasets (Figure 4).
-When the model was provided with two illustrative chest radiographs and reports, there was a marked improvement in the G&R+/R+ on the NIH dataset, with rates rising to 5.72/45 (12.7%) with a SD of 0.19, a SE of 0.03, and an IQR of 0.25.
-G&R+/G+ also enhanced to 4.69/45 (10.4%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.22; and the F1 score reached 5.03/45 (11.1%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.22.
-On the MIDRC dataset, the G&R+/R+ was improved to 16.15/45 (35.9%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50; the G&R+/G+ improved to 16.68/45 (37.1%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50; and the F1 score improved to 15.47/45 (34.3%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50.
+When the model was provided with two illustrative chest radiographs and reports, there was a marked improvement in the average PPV on the NIH dataset, with rates rising to 5.72/45 (12.7%) with a SD of 0.19, a SE of 0.03, and an IQR of 0.25.
+Average TPR also enhanced to 4.69/45 (10.4%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.22; and the average F1 score reached 5.03/45 (11.1%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.22.
+On the MIDRC dataset, the average PPV was improved to 16.15/45 (35.9%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50; the average TPR improved to 16.68/45 (37.1%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50; and the average F1 score improved to 15.47/45 (34.3%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50.
 When tasked with detecting both ICD-10 codes and their corresponding lateralities, the GPT-4V model demonstrated improved efficacy.
-On the NIH dataset, it displayed a G&R+/R+ of 1.62/45(3.5%), with a SD of 0.10, a SE of 0.01, and an IQR of 0.0; a G&R+/G+ of 1.14/45 (2.5%), with a SD of 0.07, a SE of 0.01, and an IQR of 0.0; and an F1 score of 1.30/45(2.8%), with a SD of 0.08, a SE of 0.01, and an IQR of 0.0.
-On the MIDRC dataset, it achieved a G&R+/R+ of 8.96/45 (19.9%), with a SD of 0.22, a SE of 0.03, and an IQR of 0.33; a G&R+/G+ of 9.14/45 (20.3%), with a SD of 0.25, a SE of 0.04, and an IQR of 0.33; and an F1 score of 8.53/45 (19.0%), with a SD of 0.21, a SE of 0.03, and an IQR of 0.31.
+On the NIH dataset, it displayed an average PPV of 1.62/45(3.5%), with a SD of 0.10, a SE of 0.01, and an IQR of 0.0; an average TPR of 1.14/45 (2.5%), with a SD of 0.07, a SE of 0.01, and an IQR of 0.0; and an average F1 score of 1.30/45(2.8%), with a SD of 0.08, a SE of 0.01, and an IQR of 0.0.
+On the MIDRC dataset, it achieved an average PPV of 8.96/45 (19.9%), with a SD of 0.22, a SE of 0.03, and an IQR of 0.33; an average TPR of 9.14/45 (20.3%), with a SD of 0.25, a SE of 0.04, and an IQR of 0.33; and an average F1 score of 8.53/45 (19.0%), with a SD of 0.21, a SE of 0.03, and an IQR of 0.31.
 Figure 4: Performance of GPT-4 with Vision (GPT-4V) in the detection of radiological findings from chest radiography in the few-shot setting, with statistical significance assessed using two-tailed t-test.
@@
 b, Detecting both the radiological findings in ICD codes and their corresponding lateralities.
-* - p < 0.05; ** - p < 0.01; *** - p < 0.001; **** - p < 0.0001; ns – Not significant.
 ICD= International Classification of Diseases, Tenth Revision.
-When contrasting zero-shot and few-shot learning approaches on both datasets (Figure 5), there were improved G&R+/G+ and F1 scores in few-shot learning in both scenarios (ICD-10 codes only and ICD-10 codes with laterality).
-Nonetheless, there was not a substantial increase in the G&R+/R+ suggesting that while few-shot learning may enhance the model’s capacity to detect findings, it does not noticeably enhance the precision of the predictions.
+When contrasting zero-shot and few-shot learning approaches on both datasets (Figure 5), there were improved average TPR and F1 scores in few-shot learning in both scenarios (ICD-10 codes only and ICD-10 codes with laterality).
+Nonetheless, there was not a substantial increase in the average PPV suggesting that while few-shot learning may enhance the model’s capacity to detect findings, it does not noticeably enhance the precision of the predictions.
 Figure 5: Difference of performance of GPT-4 with Vision (GPT-V4) in the detection of radiological findings from chest radiography between zero-shot and few-shot settings, with statistical significance assessed using two-tailed t-test.
@@
 This study assessed the feasibility of using GPT-4V to detect radiological findings from chest radiographs in both zero-shot and few-shot learning contexts.
-The results (G&R+/R+ of 5.53/45 (12.3%) and G&R+/G+ of 2.60/45 (5.8%)) demonstrated that radiological findings tables generated by the GPT-4V model still need to be prepared for use in clinical practice.
+The results (average PPV of 5.53/45 (12.3%) and average TPR of 2.60/45 (5.8%)) demonstrated that radiological findings tables generated by the GPT-4V model still need to be prepared for use in clinical practice.
 We acknowledged a limitation in employing GPT-4 for converting radiological reports into a structured table, where inaccuracies in ICD-10 code assignments and distinguishing between radiological findings and conclusions may impact the data’s reliability and interpretability.
@@
 This will involve including multiple modalities, such as brain CT scans and MRIs, to conduct a more thorough evaluation of GPT-4V’s performance.
+Acknowledgment
+This work was supported by the National Institutes of Health under Award No. 4R00LM013001 (Peng), NSF CAREER Award No. 2145640 (Peng), and Amazon Research Award (Peng, Shih).
+We acknowledge parts of this article were generated with GPT-4V (powered by OpenAI’s language model; https://openai.com).

### v3

@@
-Evaluating GPT-V4 (GPT-4 with Vision) on Detection of Radiological Findings on Chest Radiographs
+Evaluating GPT-4V (GPT-4 with Vision) on Detection of Radiologic Findings on Chest Radiographs
 Abstract
 Background
-Generating radiological findings from chest radiographs is pivotal in medical image analysis.
-The emergence of GPT-4 with vision (GPT-4V) has opened new perspectives on the potential for automated image-text pair generation.
+Generating radiologic findings from chest radiographs is pivotal in medical image analysis.
+The emergence of OpenAI’s generative pretrained transformer, GPT-4 with vision (GPT-4V)[1], has opened new perspectives on the potential for automated image-text pair generation.
 However, the application of GPT-4V to real-world chest radiography is yet to be thoroughly examined.
 Purpose
-To investigate GPT-4V’s capability to generate radiological findings from real-world chest radiographs.
+To investigate GPT-4V’s capability to generate radiologic findings from real-world chest radiographs.
 Materials and Methods
-In this retrospective study, 100 chest radiographs with free-text radiology reports were annotated by a cohort of two radiology-attending physicians and three residents to establish a reference standard.
+In this retrospective study, 100 chest radiographs with free-text radiology reports were annotated by a cohort of radiologists, two attending physicians and three residents, to establish a reference standard.
 Out of 100 chest radiographs, 50 were randomly picked from the National Institutes of Health (NIH) chest radiography data set, and 50 were randomly selected from the Medical Imaging and Data Resource Center (MIDRC).
 The performance of GPT-4V at detecting imaging findings from each chest radiograph was assessed in the zero-shot settings (where it operates without prior examples) and few-shot settings (where it operates with two examples).
-Its outcomes were compared with the reference standard about radiological conditions, their corresponding codes in the International Classification of Diseases, Tenth Revision (ICD-10 codes), and laterality.
+Its outcomes were compared with the reference standard about clinical conditions and their corresponding codes in the International Classification of Diseases, Tenth Revision (ICD-10 codes), including the anatomic location (hereafter, laterality).
 Results
@@
 When both the ICD-10 codes and their corresponding laterality were considered, GPT-4V produced an average PPV of 7.8%, an average TPR of 3.5%, and an average F1 score of 4.5% on the NIH data set, and an average PPV of 3.6%, an average TPR of 4.9%, and an average F1 of 6.4% on the MIDRC data set.
-In few-shot learning, GPT-4V showed improved performance on both datasets.
-When contrasting zero-shot and few-shot learning, there were improved average TPR and F1 scores in few-shot learning.
+With few-shot learning, GPT-4V showed improved performance on both data sets.
+When contrasting zero-shot and few-shot learning, there were improved average TPRs and F1 scores in few-shot setting.
 Nonetheless, there was not a substantial increase in the average PPV.
 Conclusion
-Although GPT-4V showed promise in understanding real-world images but had limited effectiveness in interpreting real-world chest radiographs.
+Although GPT-4V has shown promise in understanding nature images, it had limited effectiveness in interpreting real-world chest radiographs.
 1 Introduction
-Generating radiological findings from chest radiographs is pivotal in medical image analysis[1].
+Generating radiologic findings from chest radiographs is pivotal in medical image analysis[2].
 Recent advancements in fine-tuned pretrained models have showcased their capability to translate image content into text[3].
-However, these models are often trained on extensive, non-specific datasets and may need more domain-specific tuning for chest radiographs.
-The emergence of OpenAI’s GPT-4 with vision (GPT-4V), a multimodal large language model (LLM) with visual recognition, has opened new perspectives on the potential for automated image-text pair generation in the medical care domain.
+However, these models are often trained on extensive nonspecific data sets and may need more domain-specific tuning for chest radiographs.
+The emergence of OpenAI’s generative pretrained transformer, GPT-4 with vision (GPT-4V), a multimodal large language model (LLM) with visual recognition, has opened new perspectives on the potential for automated image-text pair generation in the medical care domain.
 Advanced multimodal LLMs, such as GPT-4V, can understand both text and images.
-While several studies have investigated the performance of GPT-4 in generating radiological impressions[3] and summarizing clinical trials[4], the practical application of multimodal LLMs to real-world chest radiographs is yet to be thoroughly examined.
-Motivated by this gap, this study aimed to investigate GPT-4V’s capability to generate radiological findings from real-world chest radiographs.
+While several studies have investigated the performance of GPT-4 in generating radiologic impressions[4] and summarizing clinical trials[5], the practical application of multimodal LLMs to real-world chest radiographs is yet to be thoroughly examined.
+Motivated by this knowledge gap, this study aimed to investigate the capability of GPT-4V to generate radiologic findings from real-world chest radiographs.
 2 Materials & Methods
@@
 2.1 Study Design and Data Collection
-In this retrospective study, a total of 100 chest radiographs and reports were independently annotated by a cohort of two radiology-attending physicians and three residents to establish a reference standard (Figure 1).
-Out of 100 chest radiographs, 50 chest radiographs were randomly picked from the National Institutes of Health (NIH) chest radiography data set, and their corresponding reports were dictated by one radiology-attending physician and three radiology residents[3].
+In this retrospective study, a total of 100 chest radiographs and radiology reports were independently annotated by a cohort of radiologists that included two attending physicians and three residents to establish a reference standard (Fig 1).
+Of 100 chest radiographs, 50 were randomly selected from the National Institutes of Health (NIH) chest radiography data set, and their corresponding reports were dictated by one radiology attending physician and three radiology residents[4].
 These 50 patients have been previously reported[5].
-The prior article dealt with the generation of the impression section using the finding section in the reports, whereas this article reports on the generation of a radiological finding table from the images.
+The prior article dealt with the generation of the impression section by using the findings section in the report, whereas the current article describes the generation of a table of radiologic finding from the image.
 The remaining 50 chest radiographs and de-identified free-text radiology reports were randomly selected from the Medical Imaging and Data Resource Center (MIDRC)[5].
 Each report included a findings section and an impressions section.
-Out of 100 chest radiographs and reports, 10 cases were randomly selected (5 from the NIH dataset and 5 from the MIDRC dataset), and two were randomly selected to serve as few-shot examples for the GPT-4V model.
-The remaining 90 were used to evaluate GPT-4V’s performance in zero-shot learning scenarios, where it operates without prior examples, and in few-shot learning scenarios, where it operates with two examples.
-Its outcomes were then compared with the reference standard annotated by radiologists about clinical conditions, their corresponding codes in the International Classification of Diseases, Tenth Revision (ICD-10), and laterality.
-Figure 1: The workflow of data construction and model application. GPT-4V=GPT-4 with vision, MIDRC= Medical Imaging and Data Resource Center, NIH= National Institutes of Health
+Out of 100 chest radiographs and reports, 10 cases were randomly selected (five from the NIH data set and five from the MIDRC data set), and two were randomly selected to serve as few-shot examples for the GPT-4V model.
+The remaining 90 were used to evaluate the performance of GPT-4V in zero-shot learning setting, where it operates without prior examples, and in few-shot learning setting, where it operates with two examples.
+Its outcomes were then compared with the reference standard annotated by radiologists with regard to clinical conditions, their corresponding codes in the International Classification of Diseases, Tenth Revision (ICD-10), including laterality.
+Figure 1: Diagram shows the study workflow, including construction of data and application of GPT-4 and GPT-4 with vision (GPT-4V).
+CXR= chest radiograph, MIDRC= Medical Imaging and Data Resource Center, NIH= National Institutes of Health
 2.2 GPT-4 with Vision (GPT-4V)
-GPT-4V (Oct.13th, 2023 version; OpenAI) was used. This version of GPT enables users to instruct GPT-4 to analyze image inputs.
+GPT-4V (Oct.13th, 2023 version; OpenAI) was used in this study[1]. GPT-4V is a version of GPT-4 that allows users to instruct the LLM to analyze image inputs.
 2.3 Experimental Setup
-To obtain the reference standard tables, GPT-4 was used to convert each free-text radiology report (Figure 2a) into a table of radiological findings with the prompt shown in Figure 2b.
-This table included the radiological findings, the corresponding ICD-10 diagnostic codes and their laterality, and detailed descriptions of the ICD codes (Figure 2c).
+To obtain the reference standard tables, GPT-4 was used to convert each free-text radiology report (Fig 2A) into a table of radiologic findings by using a textual prompt (Appendix A).
+This table includes the radiologic findings, the corresponding ICD-10 diagnostic codes, and their laterality, as well as detailed descriptions of the ICD codes (Table S1).
 Subsequently, each report was independently evaluated by three readers from a cohort of five board-certified radiologists and residents (H.O., P.K., C.W., J.K., G.S).
-Two of the readers were third-year radiology residents, and the remaining three are radiology attendings, each with over 15 years of experience.
-Their radiology subspecialties cover chest, ED/bone, neurology, and body imaging.
+Two of the readers were 3rd-year radiology residents and the remaining three are radiology attending physicians, each with over 15 years of experience.
+Their radiology subspecialties cover chest, emergency department, bone, neurology, and body imaging.
 All readers had access to the image views and reports but not to additional clinical or patient data.
-Both datasets, the National Institutes of Health (NIH) and the Medical Imaging and Data Resource Center (MIDRC), were comprehensively reviewed, with each dataset’s 50 reports being examined by three reviewers to maintain consistency and objectivity in the evaluation process.
-Findings were only included in the final tables if they were observed by at least two of the three reviewers; conversely, findings were excluded if two or more reviewers did not identify them.
-The majority vote principle was employed to provide a clear consensus for the presence or absence of radiological findings and lead to the final reference-standard table.
-Figure 2: An example.
-a, The radiology report.
-b, The prompt used to convert the free-text report into a table of radiological findings.
-c, The table of radiological findings that has been confirmed by radiologists using the report.
-d, The chest radiography input of GPT-V.
-e, The prompt utilized by GPT-4V to create a table of radiological findings derived from the chest radiography.
-f, The result produced by GPT-4V.
+Both data sets, the National Institutes of Health (NIH) and the Medical Imaging and Data Resource Center (MIDRC), were comprehensively reviewed, with the 50 reports of each data set being examined by three readers to maintain consistency and objectivity in the evaluation process.
+Findings were only included in the final tables if they were observed by at least two of the three readers; conversely, findings were excluded if two or more readers did not identify them.
+The majority vote principle was employed to provide a clear consensus for the presence or absence of radiologic findings and lead to the final reference standard table for each radiograph.
+Figure 2: An example of GPT-4 with vision (GPT-4V) inputs and output, including the (A) radiology report, (B) chest radiograph, (C) prompt provided to GPT-4V to create a table of radiologic findings derived from the chest radiograph, and (D) resultant table of radiographic findings generated by GPT-4V.
+AP = anteroposterior.
 2.3.1 Evaluating Performance in the Zero-Shot Setting
-To evaluate GPT-4V’s performance in the zero-shot setting, the chest radiographs (Figure 2d) were input into the GPT-4V model (version from Oct.13th)[2] along with the prompt (Figure 2e).
-This step aims to generate a radiological findings table (Figure 2f) comparable to the reference standard.
-The analysis concentrated on aligning positive radiological finding identification and laterality between GPT-generated and consensus tables.
-Positive predictive value (PPV), true positive rate (TPR), and F1 scores were calculated at the report level (detailed in the section of “Evaluation metrics”).
-Notably, the application of the F1 scores represented an unconventional approach due to the potential involvement of multiple diagnoses.
+To evaluate performance of GPT-4V in the zero-shot setting, the chest radiographs (Fig 2B) were input into the GPT-4V model (accessed Oct.13th, 2023)[1][3] along with the prompt (Fig 2C).
+The aim of this step was to generate a radiologic findings table (Fig 2D) comparable to the reference standard.
+The analysis concentrated on aligning positive radiologic finding identification and laterality between GPT-generated tables and the consensus tables.
+The positive predictive value (PPV), true positive rate (TPR), and F1 score were calculated at the report level (see Statistical Analysis).
+Notably, the per-report average F1 scores used in this study as each case could have multiple diagnoses.
 These metrics were used to assess the accuracy of GPT-4V in detecting the ICD-10 codes and their respective laterality.
 2.3.2 Evaluating Performance in the Few-Shot Setting
-To evaluate GPT-4V’s performance in the few-shot setting, the input was extended to include two examples of chest radiographs with their corresponding radiological findings tables before the prompt.
+To evaluate the performance of GPT-4V in the few-shot setting, the input was extended to include two examples of chest radiographs with their corresponding radiologic findings tables before the prompt.
 From the pool of 10 chest radiographs, two were randomly selected to serve as few-shot examples for the GPT-4V model.
-Supplying the model with these examples helped to provide context, boosting the model’s capacity to generate an accurate radiological findings table.
-The same performance metrics— PPV, TPR, and F1 score— were used at the report level to assess the effectiveness of few-shot learning.
+Supplying the model with these examples helped to provide context, boosting the model’s capacity to generate an accurate radiologic findings table.
+The same performance metrics (PPV, TPR, F1 score) were used at the report level to assess the effectiveness of few-shot learning.
 2.4 Statistical Analysis
-When obtaining the final reference standard tables, the interrater agreement was reported using Cohen’s Kappa score[5].
-First, the GPT-4V was employed to detect radiological findings from each chest radiography and compared the predicted findings with those obtained from the radiologists.
-To convey the performance evaluation, positive predictive value (PPV) was utilized to denote the proportion of ICD-10 Codes correctly predicted by GPT-4V, while the true positive rate (TPR) represented the ration of true positive predictions to the total number of ICD-10 Codes identified by GPT-4V.
-Evaluation metrics of PPV, TPR, and F1 score were used to assess GPT-4V’s performance in detecting imaging findings from each chest radiograph [6].
-These predicted findings were then compared with those obtained from the radiologists.
-$F1=2*\frac{PPV*TPR}{PPV+TPR}$
-In the evaluation of detecting ICD-10 codes, a radiological finding was considered a true positive if its ICD-10 code aligned with that in the reference standard table.
-In evaluating both the radiological findings in ICD codes and their corresponding lateralities, a radiological finding was considered a true positive if both its ICD-10 code and laterality matched those in the reference standard table.
-For example, in Figure 2, while evaluating ICD-10 codes alone, there were 2 ICD-10 Codes correctly predicted by GPT-4V (J90 and J18.9), five findings predicted by GPT-4V, and four in the reference standard table.
-Therefore, PPV was 0.4 (=2/(2+3)) , TPR was 0.5 (=2/(2+2)), and F1 score was 0.44.
+When obtaining the final reference standard tables, the interrater agreement was assessed using Cohen’s Kappa coefficient[6].
+First, the GPT-4V was employed to detect radiologic findings from each chest radiography and these results were compared with the predicted findings obtained from the radiologists.
+To convey the performance evaluation, PPV was used to denote the proportion of ICD-10 codes correctly predicted by GPT-4V, while the TPR represented the ratio of true-positive predictions to the total number of ICD-10 codes identified by GPT-4V.
+PPV, TPR, and F1 score were the metrics used to assess the performance of GPT-4V in detecting imaging findings from each chest radiograph.
+The F1 socre is the harmonic mean of the PPV and TPR.
+These predicted findings were then compared with those obtained from the radiologists using the following equation:
+$F1=2\times\frac{PPV\times TPR}{PPV+TPR}$ (1)
+In the evaluation of detecting ICD-10 codes, a radiologic finding was considered a true-positive if its ICD-10 code aligned with that in the reference standard table.
+In evaluating both the radiologic findings in ICD-10 codes and their corresponding lateralities, a radiologic finding was considered a true positive if both its ICD-10 code and laterality matched those in the reference standard table.
+For example, when evaluating ICD-10 codes alone, there were two ICD-10 Codes correctly predicted by GPT-4V (J90 and J18.9), with five findings predicted by GPT-4V (Fig 2), and four in the reference standard table (Table S1).
+Therefore, PPV was 0.4 ($=\frac{2}{2+3}$) , TPR was 0.5 ($=\frac{2}{2+2}$), and F1 score was 0.44.
 After obtaining the PPV, TPR, and F1 score for each chest radiography, the macro averages were calculated for PPV, TPR, and F1 score across all 90 chest radiographs.
-In this study, the statistical significance was set at 0.05, and two-tailed t-tests were used for calculating the P values for the performance metrics of GPT-4V in detecting radiological findings from chest radiography, specifically assessing the detection of findings represented by ICD codes and their associated lateralities in both zero-shot and few-shot settings.
+[…]
 3 Results
 3.1 Performance in the Zero-Shot Setting
-The performance of the GPT-4V model in the zero-shot setting varied across the NIH and MIDRC datasets and under different circumstances (Figure 3).
-).
-In the task of detecting ICD-10 codes alone, the model attained an average PPV of 5.53/45 (12.3%), with a standard deviation (SD) of 0.25, a standard error (SE) of 0.04, and an interquartile range (IQR) of 0.20; an average TPR of 2.60/45 (5.8%), with a SD of 0.10, a SE of 0.02, and an IQR of 0.10; and an average F1 score of 3.30/45 (7.3%), with a SD of 0.13, a SE of 0.02, and an IQR of 0.14, on the NIH dataset.
-Conversely, on the MIDRC dataset, the model managed an average PPV of 11.25/45 (25.0%), with a SD of 0.24, a SE of 0.04, and an IQR of 0.33; an average TPR of 7.56/45 (16.8%), with a SD of 0.20, a SE of 0.03, and an IQR of 0.25; and an average F1 score of 8.20/45 (18.2%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.29.
-The notable differences in measurements between the two datasets were primarily due to the MIDRC dataset having fewer missing GPT-4V-generated ICD-10 codes than the NIH dataset.
-On the MIDRC dataset, GPT-4V generated 144 radiological findings, while the reference standard comprised 261 findings.
-However, on the NIH dataset, GPT-4V produced 102 radiological findings, while the reference standard comprised 220 findings.
-Nevertheless, when both the ICD-10 codes and their corresponding laterality were taken into account, the GPT-4V model in the zero-shot setting produced an average PPV of 3.5/45 (7.8%), with a SD of 0.20, a SE of 0.03, and an IQR of 0.0; an average TPR of 1.56/45 (3.5%), with a SD of 0.01, a SE of 0.01, and an IQR of 0.0; and an average F1 score of 2.05/45 (4.5%) with a SD of 0.10, a SE of 0.02, and an IQR of 0.0 on the NIH dataset; and an average PPV of 4.92/45 (3.6%), with a SD of 0.21, a SE of 0.03, and an IQR of 0.25; an average TPR of 2.19/45 (4.9%), with a SD of 0.09, a SE of 0.01, and an IQR of 0.10; and an average F1 of 2.90/45 (6.4%), with a SD of 0.11, a SE of 0.02, and an IQR of 0.15, on the MIDRC dataset.
-Figure 3: Performance of GPT-4 with Vision (GPT-4V) in the detection of radiological findings from chest radiography in the zero-shot setting, with statistical significance assessed using a two-tailed t-test.
-a, Detecting the radiological findings in ICD codes.
-b, Detecting both the radiological findings in ICD codes and their corresponding lateralities.
-ICD= International Classification of Diseases, Tenth Revision.
+The performance of the GPT-4V model in the zero-shot setting varied across the NIH and MIDRC data sets and for different circumstances (Fig 3).
+In the task of detecting ICD-10 codes alone, the model attained an average PPV of 5.53 of 45 remaining radiographs(12.3%) (SD [0.25], SE [0.04], IQR [0.20]), average TPR of 2.60 of 45 (5.8%) (SD [0.10], SE [0.02], IQR [0.10]), and average F1 score of 3.30 of 45 (7.3%) (SD [0.13], SE [0.02], IQR [0.14]) on the NIH data set.
+Conversely, on the MIDRC data set, the model managed an average PPV of 11.25 of 45 (25.0%) (SD [0.24], SE [0.04], IQR [0.33]), average TPR of 7.56 of 45 (16.8%) (SD [0.20], SE [0.03], IQR [0.25], and average F1 score of 8.20 of 45 (18.2%) (SD [0.17], SE [0.03], IQR [0.29]).
+The notable differences in measurements between the two data sets were primarily due to the MIDRC data set having fewer missing GPT-4V–generated ICD-10 codes than the NIH data set.
+On the MIDRC data set, GPT-4V generated 144 radiologic findings, while the reference standard comprised 261 findings.
+However, on the NIH data set, GPT-4V produced 102 radiologic findings, while the reference standard comprised 220 findings.
+Nevertheless, when both the ICD-10 codes and their corresponding laterality were taken into account, the GPT-4V model in the zero-shot setting produced an average PPV of 3.5 of 45 (7.8%) (SD [0.20], SE [0.03], IQR [0.0], average TPR of 1.56 of 45 (3.5%) (SD [0.01], SE [0.01], IQR [0.0]), and F1 score of 2.05 of 45 (4.5%) (SD [0.10], SE [0.02], IQR [0.0]) on the NIH data set; and an average PPV of 4.92 of 45 (10.9%) (SD [0.21], SE [0.03], IQR [0.25]), average TPR of 2.19 of 45 (4.9%) (SD [0.09], SE [0.01], IQR [0.10]), and average F1 of 2.90 of 45 (6.4%) (SD [0.11], SE [0.02], IQR [0.15]) on the MIDRC data set.
+Figure 3: Bar graphs show the performance of GPT-4 with vision (GPT-4V) in the detection of radiologic findings from chest radiographs in the zero-shot setting, with statistical significance assessed using the two-tailed t-test, according to (A) the radiologic findings in International Statistical Classification of Diseases, Tenth Revision (ICD-10) codes only and (B) both the radiologic findings in ICD-10 codes and their corresponding lateralities.
 3.2 Performance in the Few-Shot Setting
-In few-shot learning, GPT-4V showed improved performance on both NIH and MIDRC datasets (Figure 4).
-When the model was provided with two illustrative chest radiographs and reports, there was a marked improvement in the average PPV on the NIH dataset, with rates rising to 5.72/45 (12.7%) with a SD of 0.19, a SE of 0.03, and an IQR of 0.25.
-Average TPR also enhanced to 4.69/45 (10.4%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.22; and the average F1 score reached 5.03/45 (11.1%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.22.
-On the MIDRC dataset, the average PPV was improved to 16.15/45 (35.9%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50; the average TPR improved to 16.68/45 (37.1%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50; and the average F1 score improved to 15.47/45 (34.3%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50.
+With few-shot learning, GPT-4V showed improved performance on both NIH and MIDRC data sets (Fig 4).
+When the model was provided with two illustrative chest radiographs and their corresponding radiologic finding tables, there was a marked improvement in the average PPV on the NIH data set, with rates rising to 5.72/45 remaining radiographs (12.7%) (SD [0.19], SE [0.03], IQR [0.25]).
+The average TPR also enhanced to 4.69 of 45 (10.4%) (SD [0.17], SE [0.03], IQR [0.22]), while the average F1 score reached 5.03 of 45 (11.1%) (SD [0.17], SE [0.03], IQR [0.22]).
+On the MIDRC data set, the average PPV improved to 16.15 of 45 (35.9%) (SD [0.23], SE [0.03], IQR [0.50]), the average TPR improved to 16.68 of 45 (37.1%) (SD [0.23], SE [0.03], IQR [0.50]), and the average F1 score improved to 15.47 of 45 (34.3%) (SD [0.23], SE [0.03], IQR [0.50]).
 When tasked with detecting both ICD-10 codes and their corresponding lateralities, the GPT-4V model demonstrated improved efficacy.
-On the NIH dataset, it displayed an average PPV of 1.62/45(3.5%), with a SD of 0.10, a SE of 0.01, and an IQR of 0.0; an average TPR of 1.14/45 (2.5%), with a SD of 0.07, a SE of 0.01, and an IQR of 0.0; and an average F1 score of 1.30/45(2.8%), with a SD of 0.08, a SE of 0.01, and an IQR of 0.0.
-On the MIDRC dataset, it achieved an average PPV of 8.96/45 (19.9%), with a SD of 0.22, a SE of 0.03, and an IQR of 0.33; an average TPR of 9.14/45 (20.3%), with a SD of 0.25, a SE of 0.04, and an IQR of 0.33; and an average F1 score of 8.53/45 (19.0%), with a SD of 0.21, a SE of 0.03, and an IQR of 0.31.
-Figure 4: Performance of GPT-4 with Vision (GPT-4V) in the detection of radiological findings from chest radiography in the few-shot setting, with statistical significance assessed using two-tailed t-test.
-a, Detecting the radiological findings in ICD codes.
-b, Detecting both the radiological findings in ICD codes and their corresponding lateralities.
-ICD= International Classification of Diseases, Tenth Revision.
-When contrasting zero-shot and few-shot learning approaches on both datasets (Figure 5), there were improved average TPR and F1 scores in few-shot learning in both scenarios (ICD-10 codes only and ICD-10 codes with laterality).
-Nonetheless, there was not a substantial increase in the average PPV suggesting that while few-shot learning may enhance the model’s capacity to detect findings, it does not noticeably enhance the precision of the predictions.
-Figure 5: Difference of performance of GPT-4 with Vision (GPT-V4) in the detection of radiological findings from chest radiography between zero-shot and few-shot settings, with statistical significance assessed using two-tailed t-test.
-a, Detecting the radiological findings in ICD codes.
-b, Detecting both the radiological findings in ICD codes and their corresponding lateralities.
+On the NIH data set, it displayed an average PPV of 1.62 of 45 (3.5%) (SD [0.10], SE [0.01], IQR [0.0]), average TPR of 1.14 of 45 (2.5%) (SD [0.07], SE [0.01], IQR [0.0]), and average F1 score of 1.30 of 45 (2.8%) (SD [0.08], SE [0.01], IQR [0.0]).
+On the MIDRC data set, GPT-4V achieved an average PPV of 8.96 of 45 (19.9%) (SD [0.22], SE [0.03], IQR [0.33]), average TPR of 9.14 of 45 (20.3%) (SD [0.25], SE [0.04], IQR [0.33]), and average F1 score of 8.53 of 45 (19.0%) (SD [0.21], SE [0.03], IQR [0.31]).
+Figure 4: Bar graphs show the performance of GPT-4 with vision (GPT-4V) in the detection of radiologic findings from chest radiographs in the few-shot setting, with statistical significance assessed using the two-tailed t test, according to (A) the radiologic findings in International Statistical Classification of Diseases, Tenth Revision (ICD-10) codes only and (B) both the radiologic findings in ICD-10 codes and their corresponding lateralities.
+When contrasting zero-shot and few-shot learning approaches on both data sets (Fig 5), there were improved average TPR and F1 scores with few-shot learning in both scenarios (ICD-10 codes only and ICD-10 codes with laterality).
+Nonetheless, there was not a substantial increase in the average PPV, suggesting that while few-shot learning may enhance the model’s capacity to detect findings, it does not noticeably enhance the precision of the predictions.
+Figure 5: Bar graphs show the difference in performance of GPT-4 with vision (GPT-4V) in the detection of radiologic findings from chest radiographs between the zero-shot and few-shot settings, with statistical significance assessed using the two-tailed t-test, according to (A) the radiologic findings in International Statistical Classification of Diseases, Tenth Revision (ICD-10) codes only and (B) both the radiologic findings in ICD-10 codes and their corresponding lateralities.
 3.3 Inter-rater Agreement
 The study compared each reader’s outcomes to the reference standard to evaluate their interpretations (Table 1).
-On the NIH dataset, the IRA for Reviewer A was 0.78, for Reviewer C was -0.06, and for Reviewer D was 0.96.
-On the MIDRC dataset, the IRA for Reviewer B was 0.96, for Reviewer C was 0.82, and for Reviewer D was 0.99.
-Table 1: Reviewer Agreement (Cohen’s Kappa) for NIH and MIDRC.
-The Inter-rater Agreement (IRA) between each reviewer and the reference standard table for both NIH and MIDRC datasets.
-$k=1$ indicates perfect agreement among reviewers and the reference standard; $k=0$ indicates that agreement is no better than chance; $k=\ -1$ indicates perfect disagreement.
-NIH Results MIDRC Results
-Reviewer A 0.78 NA
-Reviewer B NA 0.96
-Reviewer C -0.06 0.82
-Reviewer D 0.96 0.99
+On the NIH data set, the interrater agreement was 0.78 for reader A, -0.06 for reader C, 0.96 for reader D.
+On the MIDRC data set, the interrater agreement was 0.96 for reader B, 0.82 for reader C, and 0.99 for reader D.
+Table 1: Reader Agreement for NIH and MIDRC.
+Note.
+—Interrater Agreement between each reader and the reference standard table for both NIH and MIDRC data sets were accessed with Cohen $\kappa$ statistics.
+$\kappa$ values were interpreted as follows: $\kappa=1$ indicates perfect agreement among reviewers and the reference standard; $\kappa=0$ indicates that agreement is no better than chance; $\kappa=\ -1$ indicates perfect disagreement.MIDRC = Medical Imaging and Data Resource Center, NA = not applicable, NIH = National Institutes of Health.
+NIH $\kappa$ Score MIDRC $\kappa$ Score
+Reader A 0.78 NA
+Reader B NA 0.96
+Reader C -0.06 0.82
+Reader D 0.96 0.99
 4 Discussion
-The emergence of multimodal large language models (LLMs) that can understand both text and images, such as GPT-4 with Vision (GPT-4V), shows potential for automated image-text pair generation.
+The emergence of multimodal large language models (LLMs) that can understand both text and images, such as OpenAI’s GPT-4V [1], shows potential for automated image-text pair generation.
 However, applying these models to real-world data is yet to be thoroughly examined.
-This study assessed the feasibility of using GPT-4V to detect radiological findings from chest radiographs in both zero-shot and few-shot learning contexts.
-The results (average PPV of 5.53/45 (12.3%) and average TPR of 2.60/45 (5.8%)) demonstrated that radiological findings tables generated by the GPT-4V model still need to be prepared for use in clinical practice.
-We acknowledged a limitation in employing GPT-4 for converting radiological reports into a structured table, where inaccuracies in ICD-10 code assignments and distinguishing between radiological findings and conclusions may impact the data’s reliability and interpretability.
-A notable limitation of the GPT-4V output was its failure to detect several radiological conditions based on corresponding codes in the International Classification of Diseases, Tenth Revision (ICD-10).
-Overall, the top three findings that GPT-4V could not detect were “endotracheal tube”, “central venous catheter”, and “degenerative changes of osseous structures”.
-Conversely, GPT-4V most accurately detected findings such as “chest drain”, “air-space disease”, and “lung opacity”.
+This study assessed the feasibility of using GPT-4V to detect radiologic findings from chest radiographs in both zero-shot and few-shot learning contexts.
+The results (average PPV, 5.53 of 45 [12.3%]; average TPR, 2.60 of 45 [5.8%]) demonstrated that radiologic findings tables generated by GPT-4V still need further preparation for use in clinical practice.
+We acknowledge a limitation in employing GPT-4 for converting radiologic reports into a structured table, where inaccuracies in the International Statistical Classification of Diseases, Tenth Revision (ICD-10) code assignments and distinguishing between radiologic findings and conclusions may impact the reliability and interpretability of the data.
+A notable limitation of the GPT-4V output was its failure to detect several clinical conditions based on corresponding ICD-10 codes.
+Overall, the top three findings that GPT-4V could not detect were “endotracheal tube,” “central venous catheter,” and “degenerative changes of osseous structures.” Conversely, GPT-4V most accurately detected findings such as “chest drain,” “air-space disease,” and “lung opacity.”
 Although GPT-4V showed promise in understanding real-world images[7], its effectiveness in interpreting real-world chest radiographs was limited.
 Our study had limitations.
-First, the few-shot training is that it may be more prone to generating ICD-10 codes already in the provided examples, potentially reducing the diversity of ICD-10 codes generated.
+First, few-shot training is that it may be more prone to generating ICD-10 codes already in the provided examples, potentially reducing the diversity of ICD-10 codes generated.
 Second, we were unable to access other multimodal LLMs that support image inputs.
@@
 Finally, due to the limited size of the data set and relatively low level of inter-rater agreement among the radiologists, analysis by GPT-4V may have been challenging.
-In conclusion, GPT-4 with Vision (GPT-4V) showed promise in understanding natural images but had limited effectiveness in interpreting real-world chest radiographs.
+In conclusion, GPT-4V has shown promise in understanding natural images but had limited effectiveness in interpreting real-world chest radiographs.
 Our results highlight the need for additional comprehensive development and assessments prior to incorporating the GPT-4V model into clinical practice routines.
@@
 To yield robust and generalizable results, we plan to explore larger and more diverse data sets using real-world data in future studies.
-This will involve including multiple modalities, such as brain CT scans and MRIs, to conduct a more thorough evaluation of GPT-4V’s performance.
+This will involve including multiple modalities, such as brain CT scans and MRIs, to conduct a more thorough evaluation of the performance of GPT-4V.
 Acknowledgment
@@
 We acknowledge parts of this article were generated with GPT-4V (powered by OpenAI’s language model; https://openai.com).
+Appendix A Prompt to convert the free-text report into a table of radiological findings
+Read the report and look for any important clinical findings.
+Provide a summary in a table format where the positive clinical conditions are 1 and the negative clinical conditions are 0.
+Designate each condition as left side, right side, or bilateral, Provide an ICD-10 code in a separate column for positive findings only or N/A if not applicable.
+Table columns include: [Exam No., Finding No., Clinical Finding, Left Side. Right Side, Bilateral, Midline, lCD-10 Code, iCD-10 Description]
+Additional instructions:
+1. Normal findings should be excluded from each table
+2. Group similar findings together where possible for each table
+3. Create a table
+Appendix B Reference standard radiological findings
+Supplementary Table 1: Reference standard radiological findings
+Radiologic Finding Location ICD-10 ICD-10 Description
+Endotracheal tube placement Midline Z93.0 Tracheostomy status
+Pulmonary opacities Bilateral J84.9 Interstitial pulmonary disease, unspecified
+Pleural effusion C Left Side J90 Pleural effusion, not elsewhere classified
+Pneumonia Bilateral J18.9 Pneumonia, unspecified organism